OpenStudy (anonymous):

6x^2+11x+4? Can you factor?

OpenStudy (anonymous):

Use the AC factoring method AC = 6*4 = 24 ; (b1 + b2) = 3 + 8 = 11. Replace in the trinomial the term (11x) by the 2 terms (3x ) and (8x), then factor.